OpenStudy (anonymous):

18. Border's Bookstore ordered 800 art books. On verifying the order, only 160 books were actually received. The percent of the order missing was

OpenStudy (anonymous):

divide the books recived by the books ordered multiply by 100, then subtract 100 by the number you get

OpenStudy (anonymous):

800-160/800 = 640/800 To get percentage just multiply by 100%, so: 4/5 x 100% = 80%